In a magic square there are 49 cells. they are filled using natural numbers from 6 to 54. what is the sum of numbers in a row
Answer by richard1234(7193)

The sum of all the numbers is 6+7+8+...+54 = 60*49/2 = 30*49 = 1470. There are seven rows, with each row having the same sum, so 1470/7 = 210.